Write da missing num dat makes da number sentence true: 400+500=600
Answer by richard1234(5390)

Your English teacher would appreciate it if you replace the slang words ("da" and "dat") with actual words.

400 + 500 = 600 + 300. Or, 400 + 500 - 300 = 600.